March 23, 1998

George F. McNulty Department of Mathematics, Computer Science, and Statistics University of South Carolina Columbia, SC 29208 Dear George,

This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new mackerel. HI!

This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new mackerel. HI!

This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new mackerel. HI!

This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new mackerel. HI!

This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new mackerel. HI!

This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new mackerel. HI!

This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new mackerel. HI!

This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new mackerel. HI!

This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new macros! This is just a littler letter litter to try out the new mackerel. HI!

Sincerely,

George F. McNulty Professor